OpenStudy (anonymous):

The table shows the ratio of caramel corn to cheddar corn in a bag. What is the constant of proportionality? Cheddar Corn Caramel Corn 10 15 20 30 30 45 Answers: 1.5 5 15 0.67

OpenStudy (anonymous):

@jigglypuff314 @KamiBug @FelicitySchimke

OpenStudy (anonymous):

ok, thanks so much! medal and fan

OpenStudy (shinalcantara):

Since you are looking for the ratio of caramel corn to cheddar corn, you'll have 15/10 = 1.5 30/20 = 1.5 45/30 = 1.5

OpenStudy (anonymous):

so then the answer would be 1.5?

OpenStudy (anonymous):

@shinalcantara

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

would it be 1.5 or 5?

OpenStudy (anonymous):

oh, ok thanks!

OpenStudy (anonymous):

yeppers cx

OpenStudy (anonymous):

@Brostep0s could you help me with another one?

OpenStudy (anonymous):

suree

OpenStudy (anonymous):

ill open a new q